OpenStudy (anonymous):

if f(x) = x^2 -3x and g(x) = (√x)-1, then f(g(x)) =

Nnesha (nnesha):

f(g(x)) can be written as f(sqrt{x}-1) so replace x with g(x) function into f(x) equation

OpenStudy (anonymous):

i did but i still dont get it unfortunatey

Nnesha (nnesha):

\[\huge\rm f(x)=x^2-3x\] replace x with sqrt{X} -1

Nnesha (nnesha):

what did you get ? can i see ur work plz

OpenStudy (anonymous):

i got ((√x)-1)^2 - 3((√)-1)

Nnesha (nnesha):

\[\huge\rm (\sqrt{x}-1)^2-3(\sqrt{x}-1)\] distribute 2nd parentheses by -3 and apply foil method for (sqrt{x}-1)^2

Nnesha (nnesha):

do you know how to distribute ?-3(Sqrt{X}-1) ?

OpenStudy (anonymous):

no i dont unfortunately. I dont know how to distribute the -3 to the sqaure roots or how to do the first set of parenthesis

Nnesha (nnesha):

apply distributive property distribute parentheses by outside number/ \[\huge\rm \color{reD}{a}(b+c) =\color{red}{a} ·b +\color{reD}{a}·c =\color{ReD}{a}b +\color{ReD}{a}c\]

Nnesha (nnesha):

\[-3(\sqrt{x}-1)\] multiply -3 times sqrt{x} and then -3 times -1

OpenStudy (anonymous):

oh okay. I got this for the second set of parenthesis: -3√x+3

OpenStudy (anonymous):

how do i foil square roots?

Nnesha (nnesha):

yes right now foil (sqrt{x} -1)^2 is same as \[\huge\rm (\sqrt{x}-1)(\sqrt{x}-1)\] multiply 2nd parentheses by by first term sqrt{x} and then by -1